OpenStudy (anonymous):

differentiate y=x^x, x>0

OpenStudy (anonymous):

take log on both sides

OpenStudy (anonymous):

log y=x log(x) apply product rule on rhs

OpenStudy (anonymous):

any doubt??

OpenStudy (anonymous):

taking log or ln?

OpenStudy (anonymous):

natural log ln.......

OpenStudy (anonymous):

then i get this answer so far.. 1/y=ln x.. then what?

OpenStudy (anonymous):

you forgot to use the chain rule on the left side...

OpenStudy (anonymous):

what chain rule, isn't it only on the right hand side?

OpenStudy (anonymous):

the first step is : ln y=x ln x right?

OpenStudy (anonymous):

use the product rule on the right...

OpenStudy (anonymous):

yes... take the derivative with respect to x on both sides.

OpenStudy (anonymous):

then i get this.. 1/y=ln x.. then what?

OpenStudy (anonymous):

you still didn't differentiate that left side properly.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\(\large [lnx]'=\frac{1}{x} \) , correct? so what is \(\large [lny]'= \) ??? use the chain rule...

OpenStudy (anonymous):

"y" is a function of x....

OpenStudy (anonymous):

omg, my calculus is terrible.. haha.. does this related to implicit diff? dy/dx(1/x)?

OpenStudy (anonymous):

well we are differentiating implicitly the relation: lny=xlnx

OpenStudy (anonymous):

here is the chain rule formula: \(\large [lnf(x)]'=\frac{1}{f(x)}\cdot f'(x)=\frac{f'(x)}{f(x)} \)

OpenStudy (anonymous):

i found the solution already... :P http://www.analyzemath.com/calculus/Differentiation/first_derivative.html tq:)

OpenStudy (anonymous):

glad u found it... as you can see from the solution you found, [lny]' = y'/y... what i was trying to tell u....